OpenStudy (anonymous):

Automobiles generally depreciate 14% per year. Hamid purchased a Smart Car for $25 000. Create an equation that represents the value of this new vehicle over time, in years. What would its value be after 20 years? How long will it take for the Smart Car to lose three-quarters of its original value?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (luigi0210):

Equation would be: \[\LARGE A=25,000(1-.14)^n \] Where n=number of years and A=total amount.

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

part b ? @Luigi0210

OpenStudy (luigi0210):

Plug in 20 as "n" into the equation and evaluate.

OpenStudy (anonymous):

i got 1224.35

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well actually it's $1224.36, but yea.

OpenStudy (anonymous):

part c??

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Solve this for n: \[\LARGE 6250=25,000(1-.14)^n \]

OpenStudy (anonymous):

what is n? @Luigi0210

OpenStudy (luigi0210):

Solve for it..

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

-10.5 ?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

is ir righttt???

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

How did you get a negative?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

It's about 9 years: http://www.wolframalpha.com/input/?i=6250%3D25%2C000%281%E2%88%92.14%29%5En

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

can you simplify them for me ?

OpenStudy (luigi0210):

Open a new question and post them separately